A point mass with mass m slides on a smooth rod. The rod rotates with a constant \omega around the origin on an x-y plane. Assume no external forces.  m 0 = at X

a) Find Lagrangian of the point mass and the equation of constraint.

b) Find Lagrange's equation of motion and eliminate the constraint.

c) Write the Lagrange's equation with undetermined multipliers and determine the constraint force.
Answer #1


u Sol a We know, Lagrangian, L = T-2 here, T = 1 mv + Imro m ✓ Joswt dr d = dt 2 rw do = 7 dt r T = 1 m (% + 20 v=0 1 mrrt mr - nw=0 &8 =w. dt -wt :r(t) = A é + Bent With undetermined multiplier. d Cal af + at Car 2L = 0 ar ar for y-ro mr-mro no=0 mlr